OpenStudy (anonymous):

FIND THE DIFERECNCE 9 4 ------ minus ----- 3x 2x^2 3x 2x2

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

i sthis \[\frac{9}{3x}-\frac{4}{2x^2}\]?

OpenStudy (anonymous):

yeah

OpenStudy (misty1212):

then lcd is \(6x^2\)

OpenStudy (misty1212):

you can do \[\frac{9\times 2x-4\times 3}{6x^2}\]

OpenStudy (anonymous):

i dont know how to solve

OpenStudy (anonymous):

\[18x -12 \over 5x ^{2}\]

OpenStudy (arbershabani97):

it's 18x-12 over 6x^2 so yoou can simplify by 6 and you get 3x-2 over ....